# AnyV2V: A Tuning-Free Framework For Any Video-to-Video Editing Tasks

### v3

@@ Abstract
In the dynamic field of digital content creation using generative models, state-of-the-art video editing models still do not offer the level of quality and control that users desire. Previous works on video editing either extended from image-based generative models in a zero-shot manner or necessitated extensive fine-tuning, which can hinder the production of fluid video edits. Furthermore, these methods frequently rely on textual input as the editing guidance, leading to ambiguities and limiting the types of edits they can perform. Recognizing these challenges, we introduce AnyV2V, a novel tuning-free paradigm designed to simplify video editing into two primary steps: (1) employing an off-the-shelf image editing model to modify the first frame, (2) utilizing an existing image-to-video generation model to generate the edited video through temporal feature injection. AnyV2V can leverage any existing image editing tools to support an extensive array of video editing tasks, including prompt-based editing, reference-based style transfer, subject-driven editing, and identity manipulation, which were unattainable by previous methods. AnyV2V can also support any video length. Our evaluation indicates that AnyV2V significantly outperforms other baseline methods in automatic and human evaluations by significant margin, maintaining visual consistency with the source video while achieving high-quality edits across all the editing tasks.
@@
1 Introduction
The development of deep generative models [23] has led to significant advancements in content creation and manipulation, especially in digital images [43, 36, 4, 28, 9, 30]. However, video generation and editing have not reached the same level of advancement as images [51, 6, 7, 24]. In the context of video editing, training a large-scale video editing model presents considerable challenges due to the scarcity of paired data and the substantial computational resources required.
To overcome these challenges, researchers proposed various approaches [16, 12, 52, 53, 32, 31, 19, 12, 60, 39, 5, 56, 27, 20], which can be categorized into two types: (1) zero-shot adaptation from pre-trained text-to-image (T2I) models or (2) fine-tuned motion module from a pre-trained T2I or text-to-video (T2V) models. The zero-shot methods [16, 12, 27, 60] usually suffer from flickering issues due to a lack of temporal understanding. On the other hand, fine-tuning methods [52, 8, 53, 19, 20] require more time and computational overhead to edit videos. Moreover, all the methods can only adhere to certain types of edits. For example, a user might want to perform edits on visuals that are out-of-distribution from the learned text encoder (e.g. change a person to a character from their artwork). Thus, a highly customizable solution is more desired in video editing applications. It would be ideal if there were methods that allowed for a seamless combination of human artistic input and the assistance provided by AI, allowing for a synergy between human effort and AI, maintaining creators’ originality and creativity while producing high-quality results.
@@ 1 Introduction
Our objective is to propagate the edited first frame across the entire video while ensuring alignment with the source video. To achieve this, we employ I2V models [59, 10, 42] for DDIM inversion to enable first-frame conditioning. With the inverted latents as initial noise and the modified first frame as the conditional signal, the I2V model can generate videos that are not only faithful to the edited first frame but also follow the appearance and motion of the source video. To further enforce the consistency of the appearance and motion with the source video, we perform feature injection in the I2V model. The two-stage editing process effectively offloads the editing operation to existing image editing tools. This design (detail in Section 3) helps AnyV2V excel in:
@@
  • Simplicity: It does not require any fine-tuning nor additional video features like previous works [19, 53, 38] to achieve high appearance and temporal consistency for video editing tasks.
From our findings, without any fine-tuning, AnyV2V can perform video editing tasks beyond the scope of current publicly available methods, such as reference-based style transfer, subject-driven editing, and identity manipulation. AnyV2V also can perform prompt-based editing and achieve superior results on common video editing evaluation metrics compared to the baseline models [16, 12, 53]. We show both quantitatively and qualitatively that our method outperforms existing SOTA baselines in Section 4.2 and Appendix B. AnyV2V is favoured in 69.7% of samples for prompt alignment and 46.2% overall preference in human evaluation, while the best baseline only achieves 31.7% and 20.7%, respectively. Our method also reaches the highest CLIP-Text score of 0.2932 in text alignment and a competitive CLIP-Image score of 0.9652 in temporal consistency.
@@ 2 Related Works
Video generation has attracted considerable attention within the field [6, 7, 37, 51, 25, 57, 21, 50, 55, 10, 3, 42, 59]. However, video manipulation also represents a significant and popular area of interest. Initial attempts, such as Tune-A-Video [53] and VideoP2P [33] involved fine-tuning a text-to-image model to achieve video editing by learning the continuous motion. The concurrent works at that time such as Pix2Video [5] and Fate-Zero [39] go for zero-shot approach, which leverages the inverted latent from a text-to-image model to retain both structural and motion information. The progressively propagates to other frames edits. Subsequent developments have enhanced the results but generally follow the two paradigms  [16, 52, 12, 56, 5, 38, 20, 19, 14, 8, 27, 60, 11]. Control-A-Video [8] and ControlVideo [60] leveraged ControlNet [58] for extra spatial guidance. TokenFlow [16] leveraged the nearest neighbor field and inverted latent to achieve temporally consistent edit. Fairy [52] followed both paradigms which they fine-tuned a text-to-image model and also cached the attention maps to propagate the frame edits. VideoSwap [19] requires additional parameter tuning and video feature extraction (e.g. tracking, point correspondence, etc) to ensure appearance and temporal consistency. CoDeF [38] allows the first image edit to propagate the other frames with one-shot tuning. UniEdit [2] leverages the inverted latent and feature maps injection to achieve a wide range of video editing with a pre-trained text-to-video model [51].
@@
Tune-A-Video [53] Stable Diffusion [43]
Pix2Video [5] SD-Depth  [43]
Gen-1 [14] Stable Diffusion [43]
TokenFlow [16] Stable Diffusion [43]
FLATTEN [12] Stable Diffusion [43]
Fairy [52] Stable Diffusion [43]
ControlVideo [60] ControlNet [58]
CoDeF [38] ControlNet [58]
VideoSwap [19] AnimateDiff [20]
UniEdit [2] Any T2V [51]
AnyV2V (Ours) Any I2V [42, 59, 10]
However, none of the methods can offer precise control to users, as the edits may not align with the user’s exact intentions or desired level of detail, often due to the ambiguity of natural language and the constraints of the model’s capabilities. For example, VideoP2P [32] is restricted to only word-swapping prompts due to the reliance on cross-attention. There is a clear need for a more precise and comprehensive solution for video editing tasks. Our work AnyV2V is the first work to empower a diverse array of video editing tasks. We compare AnyV2V with the existing methods in Table 1. As can be seen, our method excels in its applicability and compatibility.
@@ 3 AnyV2V
Our method presents a two-stage approach to video editing. Given a source video $V^{S}=\{I_{1},I_{2},I_{3},...,I_{n}\}$, where $I_{i}$ is the frame at time $i$ and $n$ denotes the video length, we first extract the initial frame $I_{1}$ and pass it into an image editing model $\phi_{\text{img}}$ to obtain an edited first frame $I^{*}_{1}=\phi_{\text{img}}(I_{1},C)$. $C$ is the auxiliary conditions for image editing models such as text prompt, mask, style, etc. In the second stage, we feed the edited first frame $I^{*}_{1}$ and a target prompt $\mathbf{s}^{*}$ into an I2V generation model $\epsilon_{\theta}$ and employ the inverted latent from the source video $V^{S}$ to guide the generation process such that the edited video $V^{*}$ follows the motion of the source video $V^{S}$, the semantic information represented in the edited first frame $I^{*}_{1}$, and the target prompt $\mathbf{s^{*}}$. An overall illustration of our video editing pipeline is shown in Figure 2. We provide a discussion of the preliminaries of our method in Appendix C. In this section, we explain each core component of our method.
@@ 3.1 Flexible First Frame Editing
In visual manipulation, controllability is a key element in performing precise editing. AnyV2V enables more controllable video editing by utilizing image editing models to modify the video’s first frame. This strategic approach enables highly accurate modifications in the video and is compatible with a broad spectrum of image editing models, including other deep learning models that can perform image style transfer [15, 17, 34, 48], mask-based image editing [36, 1], image inpainting [45, 29], identity-preserving image editing [49], and subject-driven image editing [9, 30, 18].
@@ 3.3 Appearance Guidance via Spatial Feature Injection
To enforce consistency with the source video, we perform feature injection in both convolution layers and spatial attention layers in the denoising U-Net. During the video sampling process, we simultaneously denoise the source video using the previously collected DDIM inverted latents $\mathbf{z}^{S}_{t}$ at each time step $t$ such that $\mathbf{z}^{S}_{t-1}=\epsilon_{\theta}(\mathbf{z}^{S}_{t},I_{1},\varnothing,t)$. We preserve two types of features during source video denoising: convolution features $f^{l_{1}}$ before skip connection from the $l_{1}^{\text{th}}$ residual block in the U-Net decoder, and the spatial self-attention scores $\{A_{s}^{l_{2}}\}$ from $l_{2}=\{l_{low},l_{low+1},...,l_{high}\}$ layers. We collect the queries $\{Q_{s}^{l_{2}}\}$ and keys $\{K_{s}^{l_{2}}\}$ instead of directly collecting $A_{s}^{l_{2}}$ as the attention score matrices are parameterized by the query and key vectors. We then replace the corresponding features during denoising the edited video in both the normal denoising branch and the negative prompt branch for classifier-free guidance [22]. We use two thresholds $\tau_{conv}$ and $\tau_{sa}$ to control the convolution and spatial attention injection to only happen in the first $\tau_{conv}$ and $\tau_{sa}$ steps during video sampling.
@@ 3.4 Motion Guidance through Temporal Feature Injection
The spatial feature injection mechanism described in Section 3.3 significantly enhances the background and overall structure consistency of the edited video. While it also helps maintain the source video motion to some degree, we observe that the edited videos will still have a high chance of containing incorrect motion compared to the source video. On the other hand, we notice that I2V generation models, or video diffusion models in general, are often initialized from pre-trained T2I models and continue to be trained on video data. During the training process, parameters in the spatial layers are often frozen or set to a lower learning rate such that the pre-trained weights from the T2I model are less affected, and the parameters in the temporal layers are more extensively updated during training. Therefore, it is likely that a large portion of the motion information is encoded in the temporal layers of the I2V generation models. Concurrent work [2] also observes that features in the temporal layers show similar characteristics with optical flow [26], a pattern that is often used to describe the motion of the video.
@@ 4.1 Implementation Details
We employ AnyV2V on three off-the-shelf I2V generation models: I2VGen-XL [^1] We use the version provided in https://huggingface.co/ali-vilab/i2vgen-xl. [59], ConsistI2V [42] and SEINE [10]. For all I2V models, we use $\tau_{conv}=0.2T$, $\tau_{sa}=0.2T$ and $\tau_{ta}=0.5T$, where $T$ is the total number of sampling steps. We use the DDIM [44] sampler and set $T$ to the default values of the selected I2V models. Following PnP [47], we set $l_{1}=4$ for convolution feature injection and $l_{2}=l_{3}=\{4,5,6,...,11\}$ for spatial and temporal attention injections. During sampling, we apply text classifier-free guidance (CFG) [22] for all models with the same negative prompt “Distorted, discontinuous, Ugly, blurry, low resolution, motionless, static, disfigured, disconnected limbs, Ugly faces, incomplete arms” across all edits. To obtain the initial edited frames in our implementation, we use a set of image editing model candidates including prompt-based image editing model InstructPix2Pix [4], style transfer model Neural Style Transfer (NST) [15], subject-driven image editing model AnyDoor [9], and identity-driven image editing model InstantID [49]. We experiment with only the successfully edited frames, which is crucial for our method. We conducted all the experiments on a single Nvidia A6000 GPU. To edit a 16-frame video, it requires around 15G GPU memory and around 100 seconds for the whole inference process. We refer readers to Appendix A for more discussions on our implementation details and hyperparameter settings.

### v4

@@ Abstract
In the dynamic field of digital content creation using generative models, state-of-the-art video editing models still do not offer the level of quality and control that users desire. Previous works on video editing either extended from image-based generative models in a zero-shot manner or necessitated extensive fine-tuning, which can hinder the production of fluid video edits. Furthermore, these methods frequently rely on textual input as the editing guidance, leading to ambiguities and limiting the types of edits they can perform. Recognizing these challenges, we introduce AnyV2V, a novel tuning-free paradigm designed to simplify video editing into two primary steps: (1) employing an off-the-shelf image editing model to modify the first frame, (2) utilizing an existing image-to-video generation model to generate the edited video through temporal feature injection. AnyV2V can leverage any existing image editing tools to support an extensive array of video editing tasks, including prompt-based editing, reference-based style transfer, subject-driven editing, and identity manipulation, which were unattainable by previous methods. AnyV2V can also support any video length. Our evaluation shows that AnyV2V achieved CLIP-scores comparable to other baseline methods. Furthermore, AnyV2V significantly outperformed these baselines in human evaluations, demonstrating notable improvements in visual consistency with the source video while producing high-quality edits across all editing tasks. The code is available at https://github.com/TIGER-AI-Lab/AnyV2V.
1 Introduction
The development of deep generative models (Ho et al., 2020) has led to significant advancements in content creation and manipulation, especially in digital images (Rombach et al., 2022; Nichol et al., 2022; Brooks et al., 2023; Ku et al., 2024; Chen et al., 2023c; Li et al., 2023). However, video generation and editing have not reached the same level of advancement as images (Wang et al., 2023; Chen et al., 2023a; 2024; Ho et al., 2022). In the context of video editing, training a large-scale video editing model presents considerable challenges due to the scarcity of paired data and the substantial computational resources required.
To overcome these challenges, researchers proposed various approaches (Geyer et al., 2023; Cong et al., 2023; Wu et al., 2023a; b; Liu et al., 2023a; Liang et al., 2023; Gu et al., 2023b; Cong et al., 2023; Zhang et al., 2023d; Qi et al., 2023; Ceylan et al., 2023; Yang et al., 2023; Jeong & Ye, 2023; Guo et al., 2023), which can be categorized into two types: (1) zero-shot adaptation from pre-trained text-to-image (T2I) models or (2) fine-tuned motion module from a pre-trained T2I or text-to-video (T2V) models. The zero-shot methods (Geyer et al., 2023; Cong et al., 2023; Jeong & Ye, 2023; Zhang et al., 2023d) usually suffer from flickering issues due to a lack of temporal understanding. On the other hand, fine-tuning methods (Wu et al., 2023a; Chen et al., 2023b; Wu et al., 2023b; Gu et al., 2023b; Guo et al., 2023) require more time and computational overhead to edit videos. Moreover, all the methods can only adhere to certain types of edits. For example, a user might want to perform edits on visuals that are out-of-distribution from the learned text encoder (e.g. change a person to a character from their artwork). Thus, a highly customizable solution is more desired in video editing applications. It would be ideal if there were methods that allowed for a seamless combination of human artistic input and the assistance provided by AI, allowing for a synergy between human effort and AI, maintaining creators’ creativity while producing high-quality outputs.
@@ 1 Introduction
Our objective is to propagate the edited first frame across the entire video while ensuring alignment with the source video. To achieve this, we employ I2V models (Zhang et al., 2023c; Chen et al., 2023d; Ren et al., 2024) for DDIM inversion to enable first-frame conditioning. With the inverted latents as initial noise and the modified first frame as the conditional signal, the I2V model can generate videos that are not only faithful to the edited first frame but also follow the appearance and motion of the source video. To further enforce the consistency of the appearance and motion with the source video, we perform feature injection in the I2V model. The two-stage editing process effectively offloads the editing operation to existing image editing tools. This design (detail in Section 4) helps AnyV2V excel in:
@@
  • Simplicity: It does not require any fine-tuning nor additional video features like previous works (Gu et al., 2023b; Wu et al., 2023b; Ouyang et al., 2023) to achieve high appearance and temporal consistency for video editing tasks.
From our findings, without any fine-tuning, AnyV2V can perform video editing tasks beyond the scope of current publicly available methods, such as reference-based style transfer, subject-driven editing, and identity manipulation. AnyV2V also can perform prompt-based editing and achieve superior results on common video editing evaluation metrics compared to the baseline models (Geyer et al., 2023; Cong et al., 2023; Wu et al., 2023b). We show both quantitatively and qualitatively that our method outperforms existing SOTA baselines in Section 5.3 and Appendix B. AnyV2V is favoured in 69.7% of samples for prompt alignment and 46.2% overall preference in human evaluation, while the best baseline only achieves 31.7% and 20.7%, respectively. Our method also reaches the highest CLIP-Text score of 0.2932 in text alignment and a competitive CLIP-Image score of 0.9652 in temporal consistency.
@@ 2 Related Works
Video generation has attracted considerable attention within the field (Chen et al., 2023a; 2024; OpenAI, ; Wang et al., 2023; Hong et al., 2022; Zhang et al., 2023a; Henschel et al., 2024; Wang et al., 2024c; Xing et al., 2023; Chen et al., 2023d; Bar-Tal et al., 2024; Ren et al., 2024; Zhang et al., 2023c). However, video manipulation also represents a significant and popular area of interest. Initial attempts, such as Tune-A-Video (Wu et al., 2023b) and VideoP2P (Liu et al., 2023b) involved fine-tuning a text-to-image model to achieve video editing by learning the continuous motion. The concurrent works at that time such as Pix2Video (Ceylan et al., 2023) and Fate-Zero (Qi et al., 2023) go for zero-shot approach, which leverages the inverted latent from a text-to-image model to retain both structural and motion information. The progressively propagates to other frames edits. Subsequent developments have enhanced the results but generally follow the two paradigms  (Geyer et al., 2023; Wu et al., 2023a; Cong et al., 2023; Yang et al., 2023; Ceylan et al., 2023; Ouyang et al., 2023; Guo et al., 2023; Gu et al., 2023b; Esser et al., 2023; Chen et al., 2023b; Jeong & Ye, 2023; Zhang et al., 2023d; Cheng et al., 2023). Control-A-Video (Chen et al., 2023b) and ControlVideo (Zhang et al., 2023d) leveraged ControlNet (Zhang et al., 2023b) for extra spatial guidance. TokenFlow (Geyer et al., 2023) leveraged the nearest neighbor field and inverted latent to achieve temporally consistent edit. Fairy (Wu et al., 2023a) followed both paradigms which they fine-tuned a text-to-image model and also cached the attention maps to propagate the frame edits. VideoSwap (Gu et al., 2023b) requires additional parameter tuning and video feature extraction (e.g. tracking, point correspondence, etc) to ensure appearance and temporal consistency. CoDeF (Ouyang et al., 2023) allows the first image edit to propagate the other frames with one-shot tuning. UniEdit (Bai et al., 2024) leverages the inverted latent and feature maps injection to achieve a wide range of video editing with a pre-trained text-to-video model (Wang et al., 2023).
@@
Tune-A-Video (Wu et al., 2023b) Stable Diffusion
Pix2Video (Ceylan et al., 2023) SD-Depth
Gen-1 (Esser et al., 2023) Stable Diffusion
TokenFlow (Geyer et al., 2023) Stable Diffusion
FLATTEN (Cong et al., 2023) Stable Diffusion
Fairy (Wu et al., 2023a) Stable Diffusion
ControlVideo (Zhang et al., 2023d) ControlNet
CoDeF (Ouyang et al., 2023) ControlNet
VideoSwap (Gu et al., 2023b) AnimateDiff
UniEdit (Bai et al., 2024) Any T2V Models
AnyV2V (Ours) Any I2V Models
However, none of the methods can offer precise control to users, as the edits may not align with the user’s exact intentions or desired level of detail, often due to the ambiguity of natural language and the constraints of the model’s capabilities. For example, VideoP2P (Liu et al., 2023a) is restricted to only word-swapping prompts due to the reliance on cross-attention. There is a clear need for a more precise and comprehensive solution for video editing tasks. Our work AnyV2V is the first work to empower a diverse array of video editing tasks. We compare AnyV2V with the existing methods in Table 1. As can be seen, our method excels in its applicability and compatibility.
3 Preliminary
3.1 Image-to-Video (I2V) Generation Models
In this work, we focus on leveraging latent diffusion-based (Rombach et al., 2022) I2V generation models for video editing. Given an input first frame $I_{1}$, a text prompt $\mathbf{s}$ and a noisy video latent $\mathbf{z}_{t}$ at time step $t$, I2V generation models recover a less noisy latent $\mathbf{z}_{t-1}$ using a denoising model $\epsilon_{\theta}(\mathbf{z}_{t},I_{1},\mathbf{s},t)$ conditioned on both $I_{1}$ and $\mathbf{s}$. The denoising model $\epsilon_{\theta}$ contains a set of spatial and temporal self-attention layers, where the self-attention operation can be formulated as:
$Q=W^{Q}z,K=W^{K}z,V=W^{V}z,$ (1)
3.2 DDIM Inversion
The denoising process for I2V generation models from $\mathbf{z}_{t}$ to $\mathbf{z}_{t-1}$ can be achieved using the DDIM (Song et al., 2020) sampling algorithm. The reverse process of DDIM sampling, known as DDIM inversion (Mokady et al., 2023; Dhariwal & Nichol, 2021), allows obtaining $\mathbf{z}_{t+1}$ from $\mathbf{z}_{t}$ such that $\mathbf{z}_{t+1}=\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}\mathbf{z}_{t}+(\sqrt{\frac{1}{\alpha_{t+1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1})\cdot\epsilon_{\theta}(\mathbf{z}_{t},x_{0},\mathbf{s},t)$, where $\alpha_{t}$ is derived from the variance schedule of the diffusion process.
3.3 Plug-and-Play (PnP) Diffusion Features
Tumanyan et al. (2023a) proposed PnP diffusion features for image editing, based on the observation that intermediate convolution features $f$ and self-attention scores $A=\mathrm{Softmax}(\frac{QK^{\top}}{\sqrt{d}})$ in a text-to-image (T2I) denoising U-Net capture the semantic regions (e.g. legs or torso of a human body) during the image generation process.
@@ 4 AnyV2V
Our method presents a two-stage approach to video editing. Given a source video $V^{S}=\{I_{1},I_{2},I_{3},...,I_{n}\}$, where $I_{i}$ is the frame at time $i$ and $n$ denotes the video length, we first extract the initial frame $I_{1}$ and pass it into an image editing model $\phi_{\text{img}}$ to obtain an edited first frame $I^{*}_{1}=\phi_{\text{img}}(I_{1},C)$. $C$ is the auxiliary conditions for image editing models such as text prompt, mask, style, etc. In the second stage, we feed the edited first frame $I^{*}_{1}$ and a target prompt $\mathbf{s}^{*}$ into an I2V generation model $\epsilon_{\theta}$ and employ the inverted latent from the source video $V^{S}$ to guide the generation process such that the edited video $V^{*}$ follows the motion of the source video $V^{S}$, the semantic information represented in the edited first frame $I^{*}_{1}$, and the target prompt $\mathbf{s^{*}}$. An overall illustration of our video editing pipeline is shown in Figure 2. In this section, we explain each core component of our method.
@@ 4.1 Flexible First Frame Editing
In visual manipulation, controllability is a key element in performing precise editing. AnyV2V enables more controllable video editing by utilizing image editing models to modify the video’s first frame. This strategic approach enables highly accurate modifications in the video and is compatible with a broad spectrum of image editing models, including other deep learning models that can perform image style transfer (Gatys et al., 2015; Ghiasi et al., 2017; Lötzsch et al., 2022; Wang et al., 2024a), mask-based image editing (Nichol et al., 2022; Avrahami et al., 2022), image inpainting (Suvorov et al., 2021; Ku et al., 2022), identity-preserving image editing (Wang et al., 2024b), and subject-driven image editing (Chen et al., 2023c; Li et al., 2023; Gu et al., 2023a).
@@ 4.3 Appearance Guidance via Spatial Feature Injection
To enforce consistency with the source video, we perform feature injection in both convolution layers and spatial attention layers in the denoising U-Net. During the video sampling process, we simultaneously denoise the source video using the previously collected DDIM inverted latents $\mathbf{z}^{S}_{t}$ at each time step $t$ such that $\mathbf{z}^{S}_{t-1}=\epsilon_{\theta}(\mathbf{z}^{S}_{t},I_{1},\varnothing,t)$. We preserve two types of features during source video denoising: convolution features $f^{l_{1}}$ before skip connection from the $l_{1}^{\text{th}}$ residual block in the U-Net decoder, and the spatial self-attention scores $\{A_{s}^{l_{2}}\}$ from $l_{2}=\{l_{low},l_{low+1},...,l_{high}\}$ layers. We collect the queries $\{Q_{s}^{l_{2}}\}$ and keys $\{K_{s}^{l_{2}}\}$ instead of directly collecting $A_{s}^{l_{2}}$ as the attention score matrices are parameterized by the query and key vectors. We then replace the corresponding features during denoising the edited video in both the normal denoising branch and the negative prompt branch for classifier-free guidance (Ho & Salimans, 2022). We use two thresholds $\tau_{conv}$ and $\tau_{sa}$ to control the convolution and spatial attention injection to only happen in the first $\tau_{conv}$ and $\tau_{sa}$ steps during video sampling.
@@ 4.4 Motion Guidance through Temporal Feature Injection
The spatial feature injection mechanism described in Section 4.3 significantly enhances the background and overall structure consistency of the edited video. While it also helps maintain the source video motion to some degree, we observe that the edited videos will still have a high chance of containing incorrect motion compared to the source video. On the other hand, we notice that I2V generation models, or video diffusion models in general, are often initialized from pre-trained T2I models and continue to be trained on video data. During the training process, parameters in the spatial layers are often frozen or set to a lower learning rate such that the pre-trained weights from the T2I model are less affected, and the parameters in the temporal layers are more extensively updated during training. Therefore, it is likely that a large portion of the motion information is encoded in the temporal layers of the I2V generation models. Concurrent work (Bai et al., 2024) also observes that features in the temporal layers show similar characteristics with optical flow (Horn & Schunck, 1981), a pattern that is often used to describe the motion of the video.
@@ 5.1 Implementation Details
We employ AnyV2V on three off-the-shelf I2V generation models: I2VGen-XL [^1] We use the version provided in https://huggingface.co/ali-vilab/i2vgen-xl. (Zhang et al., 2023c), ConsistI2V (Ren et al., 2024) and SEINE (Chen et al., 2023d). For all I2V models, we use $\tau_{conv}=0.2T$, $\tau_{sa}=0.2T$ and $\tau_{ta}=0.5T$, where $T$ is the total number of sampling steps. We use the DDIM (Song et al., 2020) sampler and set $T$ to the default values of the selected I2V models. Following PnP (Tumanyan et al., 2023b), we set $l_{1}=4$ for convolution feature injection and $l_{2}=l_{3}=\{4,5,6,...,11\}$ for spatial and temporal attention injections. During sampling, we apply text classifier-free guidance (CFG) (Ho & Salimans, 2022) for all models with the same negative prompt “Distorted, discontinuous, Ugly, blurry, low resolution, motionless, static, disfigured, disconnected limbs, Ugly faces, incomplete arms” across all edits. To obtain the initial edited frames in our implementation, we use a set of image editing model candidates including prompt-based image editing model InstructPix2Pix (Brooks et al., 2023), style transfer model Neural Style Transfer (NST) (Gatys et al., 2015), subject-driven image editing model AnyDoor (Chen et al., 2023c), and identity-driven image editing model InstantID (Wang et al., 2024b). We experiment with only the successfully edited frames, which is crucial for our method. We conducted all the experiments on a single Nvidia A6000 GPU. To edit a 16-frame video, it requires around 15G GPU memory and around 100 seconds for the whole inference process. We refer readers to Appendix A for more discussions on our implementation details and hyperparameter settings.